y 10 kN 50 kNm A с X +B 2m 3 m

Consider the beam and loading shown. The cross-sectional area is rectangular which measures 200 mm x 400 mm. What is Mmax in KN-m? Use two decimal places.
Answer #1

Given E= 10kn, M = 50kN-M JOEN ASOKNEM B RAH am 3m Rc From figure Some of horizontal forces forces are o BACH And sum verticaLet TEMA=0 Rx 5 -102.-50=0 R = 14 KN And RAU= 10-Rc Rau= 10-14 Bar = -14 KN BENBDING MOMENT DIAGRAM 42 KN-m tve) A B (ve) 8 k